CEFSCONV(1) CEFSCONV(1) NAME cefsconv−convertaTeXdocumentinSJISencodingcontainingCEFmacrosinto‘preprocessed’ form. SYNOPSIS cefsconv<infile>outfile DESCRIPTION TheChineseEncodingFramework(CEF)developedbyChristianWittern <cwittern@conline.central.de>makesitpossibletoincludeCJKcharactersinvariousencodingsusingtheSGMLmacros&CX-aabb;forBig5andCNSencodingand&U-aabb;forUni-code(Xdenotesanumberbetween0and7;0representsBig5encoding,1-7CNSencoding planes1-7.aabbiseitheradecimalorhexadecimalnumber). cefsconvconvertstheCEFmacrosofaSJISencodeddocumenttogetherwiththeSJISencoding itselfintoaformwhichcandirectlyprocessedwithLATEX2ε.Itisidenticalto sjisconv<infile|cefconv>outfile ThisfilterispartoftheCJKmacropackageforLATEX2ε. SEEALSO sjisconv(1),bg5conv(1),cefconv(1),cef5conv(1),extconv(1), theCJKdocumentationfiles. AUTHOR WernerLemberg<wl@gnu.org> CJKVersion4.3.0 20-Jun-1999 1